# Orders in Chaos: Enhancing Large-Scale MoE LLM Serving with Data Movement Forecasting

### v2

@@ V-A Experiment Setup
Methodology: We conduct experiments using trace-driven simulation on a validated simulator. Expert selection traces are collected by deploying SGLang [zheng2024sglang] on an 8×H100 DGX server and an 8×H200 AWS instance.
We developed a custom multi-chiplet GPU simulator in Python since existing tools are inadequate for our needs. Cycle-accurate simulators like Gem5 [gem5], gpgpusim [gpgpusim], and mgpusim [mgpusim] accurately model single GPUs but are too slow for large-scale systems with 20+ GPUs serving batch sizes exceeding 15,000. Event-driven ASTRA-sim [astra] supports multi-GPU systems but lacks MoE serving support and detailed single-GPU microarchitecture modeling. To ensure accuracy, we validated our simulator against real data from an 8×H100 DGX server, as shown in Sec. V-B.
@@
| Dojo | 5 | 5 | 2 TB/s | 1.5 TB/s | 80GB | 1000 TFLOPS | ||||||
| TSMC-SoW | 3 | 8 | 2 TB/s | 1.5 TB/s | 80GB | 1000 TFLOPS | ||||||
@@
Metric: We measure pure MoE layer performance during the decode stage as modern LLM serving systems show a trend toward fine-granularity disaggregation. Traditional LLMs benefit from separating prefill and decode stages across different machines, as demonstrated by DistServe [zhong2024distserve] and subsequent works [loongserve, qin2025mooncake]. For MoE models, this disaggregation extends further. MegaScale-Infer [zhu2025megascale-infer] separates attention and MoE operations onto different machines for optimal batch sizes. Following this trend, we focus on optimizing MoE operations during the decode stage.
Hardware Configuration: We evaluate two multi-chiplet topologies: Tesla Dojo [dojo1, dojo2] and the TSMC SoW roadmap [tsmc-arch]. As summarized in Table I, Dojo uses a 55 2D mesh, while TSMC SoW adopts an 83 2D mesh. These choices reflect a deployed system (Dojo) and near-future industry support (TSMC SoW).
For both the Dojo and TSMC SoW configurations, each chiplet is A100-like, providing 1,000 TFLOPS FP8 compute, 80GB HBM, 2TB/s local HBM bandwidth, and 1.5 TB/s inter-die bandwidth to adjacent chiplets. We also include an extended experiment in Sec. V-F with a Dojo-Enhanced configuration, where each die is B300-like to reflect an anticipated hardware performance trend in the future. We reserve 10% of DRAM for system and hardware management.
Baseline Configurations: We compare our approach against the simple strategy currently used by GPU. The Base configuration assigns equal numbers of experts to each die but ignores expert placement when allocating tasks. We implement three variants: Allo Only uses solely our task allocation strategy; Pred Only includes only the data-driven predictor; and Allo+Pred combines both techniques. These configurations evaluate the individual and combined effects of our proposed methods.
Models and Workloads: We conduct evaluations with real traces collected from Qwen3 and Deepseek V3. The traces are gathered from diverse datasets, including MMLU [mmlu], MMLU Pro [mmlu_pro], ChineseSimpleQA [he2024chinese], and LiveCodeBench [jain2024livecodebench], comprising over 24,000 requests per model.
@@ V-B Validation of Simulator
We validate our simulator using real measurements from an 8×H100 DGX server. We evaluate both single-GPU execution and two-GPU peer-to-peer (P2P) communication. For single-GPU execution, we benchmark GEMM operations for experts in the MoE layer across varying batch sizes for both DeepSeek and Qwen. For P2P communication, we measure data migration between two GPUs over payload sizes ranging from 4 KB to 4 GB. To ensure simulation fidelity, we calibrate key parameters to fit the measured data. As shown in Figure 12, the simulator’s error remains within 5% for all test cases.
@@ V-C Throughput
Comparison across models: Our Allo+Pred strategy achieves 5.3 throughput improvement on Deepseek and 3.1 on Qwen. Deepseek shows greater improvement due to its larger number of experts (256 vs. 128) and more complex expert selection patterns. Without our expert placement-aware allocation, serving Deepseek introduces significantly more inter-unit communication than Qwen. Since our task allocation strategy places computation close to data, it yields more substantial improvements on Deepseek.
Comparison across chiplet architectures: Our Allo+Pred strategy shows 3.7 improvement on Dojo and 4.7 on TSMC, despite similar die counts (25 vs. 24). The improvement on TSMC is more substantial due to its rectangular layout, where dies are farther apart, introducing more inter-unit communication without strategic task allocation.
@@ V-D Hop Reduction
We report hop counts in Figure 11 to show the reduction in inter-unit communication. Hop count is the sum of Manhattan distances for all cross-unit communications. Higher hop counts indicate frequent cross-die data movement. We normalize results to baseline and report hop reduction ratios, where a ratio of 10 means hop count is reduced to 1/10.
Pred Only reduces hop counts by 3.7-4.4, aligning with performance results in Figure 11. This indicates cross-unit communication is the primary bottleneck in baseline, and reducing hop counts proportionally improves performance.
Allo Only reduces hop counts by 7.2-10.8, exceeding the performance improvement range of 17-661. This shows that with our allocation algorithm, inter-unit communication is no longer the sole bottleneck. While reducing hop counts still improves performance, the improvement is not proportional.
Allo+Pred reduces hop counts by over 210-944 compared to baseline and achieves 10 improvement over Allo Only. However, performance improvement is only 2.5-6.5 over baseline, with just 1.2 average improvement over Allo Only. This demonstrates that hop count is no longer a performance bottleneck. Our task allocation successfully places computation near relevant data. Most tasks are distributed to local dies holding related experts, with only extremely popular experts requiring remote allocation. This leads to minimal D2D traffic and shifts the bottleneck to workload distribution.
@@ V-F Comparison with Host CPU-Based Implementation
V-G Discussion
@@
This wafer-scale GPU architecture serves as a case study demonstrating the practical applicability of our proposed insights, which constitute the paper’s primary contribution. Our design implements only a subset of the proposed insights to validate their effectiveness. We follow Insight 3 for task allocation and leverage temporal relation insights (Insight 1 and part of Insight 2) to build a data-driven predictor. However, the remaining insights offer significant potential for further exploitation in wafer-scale architectures. For instance, Insight 2 could inform LLC prefetch strategies and enhance task allocation algorithms that consider LLC data placement. Spatial relation insights (Insight 4 and Insight 5) could optimize initial expert placement, which could be further refined by incorporating task information from Insight 6. Importantly, our insights extend far beyond this case study to benefit diverse future systems, including multi-GPU clusters [zhu2025megascale-infer, zhang2025comet], CXL-based or CPU-based memory disaggregation [moelightning, fang2025klotski], Flash-based multi-hierarchy systems [cambricon-llm, sheng2023flexgen], and other emerging architectures and systems.
VI Related Work
@@
Several MoE model tech reports [muennighoff2024olmoe, jiang2024mixtral, dai2024deepseekmoe, sun2024hunyuan] provide the MoE routing patterns as part of their evaluation.
For example, the Mixtral report [jiang2024mixtral] shows the temporal locality of expert selection by reporting the percentage of repetitive assignment.
The OLMoE report [muennighoff2024olmoe] shows the co-activation pattern and domain specialization among the experts.
A blog post from SGLang [sglang-deepseek-blogpost] shows expert distribution statistics for the DeepSeekV3 model and the inherent imbalance in expert selection and similarity between prefill and decode.
@@
Various prior work [hwang2024pre, li2023accelerating, zhai2023smartmoe, gupta2024lynx, kamahori2025fiddler, eliseev2023fast, du2024sida, li2024optimizing, skliar2024mixture, moelightning, yun2024duplex, zhang2025comet, hwang2023tutel] have focused on improving the efficiency of MoE LLM inference by reducing data movement.
For example, Lina [li2023accelerating] exploits the skewness in expert selection to dynamically schedule resources during inference and balance traffic across multiple GPUs.
LYNX [gupta2024lynx] dynamically reduces active experts while preserving model accuracy.
Pre-gate MoE [hwang2024pre] uses a pre-gating function to alleviate the dynamic nature of expert selection.
Sida [du2024sida] builds an offline hash function to predict expert usage to reduce data movement between CPU and GPU.
MoE-Lightning [moelightning] leverages a CPU-GPU pipeline and paged weights to improve resource utilization.
Eliseev and Mazur [eliseev2023fast] exploit expert locality and leverage LRU caching to manage GPU and CPU memory.
This work aims to reduce data movement in MoE LLM inference of today’s systems, and our case study shows that data-driven profiling across multiple models provides a solid base to optimize not only current but future systems of any scale.
@@
As single chip scaling slows down, packaging technologies such as wafer-scale and chiplet packaging have become a promising path to further improve compute efficiency for future systems.
Recent work in this area either focuses on the interconnect design [rashidi2025fred, yang2025pd, yu2025cramming, li2024lucie, chen2024waferscale, feng2023heterogeneous, feng2023scalable] of wafer-scale and chiplet systems, or
general data placement for specific algorithms [he2025waferlm, tan2021nn] and applications [shao2019simba, xu2025wsc, odema2024scar].
In contrast, our work is the first to study the implications of serving large-scale MoE LLMs in a wafer-scale GPU and propose data-movement-centric HW/SW co-design optimizations.
VII Conclusion
In this work, unlike previous MoE serving works that adopt system-centric approaches and develop deployment-specific strategies, we study MoE model serving from a model-focused perspective We conducted comprehensive data-movement-centric profiling of the latest large-scale MoE models (200B-1000B) to extract system-agnostic insights. These insights reveal underlying patterns behind seemingly random data movement, providing actionable guidance for designing diverse future MoE serving systems. We validate our insights on a wafer-scale GPU architecture and achieve significant performance improvements with minimal modifications.

### v4

@@ V-A Experiment Setup
Methodology: We conduct experiments using event-driven simulation on a validated simulator. Expert selection traces are collected by deploying SGLang [90] on an 8×H100 DGX server and an 8×H200 AWS instance.
We developed a custom multi-chiplet GPU simulator in Python since existing tools are inadequate for our needs. Cycle-accurate simulators like Gem5 [4], gpgpusim [2], and mgpusim [62] accurately model single GPUs but are too slow for large-scale systems with 20+ GPUs serving batch sizes exceeding 15,000. Event-driven ASTRA-sim [78] supports multi-GPU systems but lacks MoE serving support and detailed single-GPU microarchitecture modeling. To ensure accuracy, we validated our simulator against real data from an 8×H100 DGX server, as shown in Sec. V-B.
@@
| Dojo | 5 | 5 | 3.35 TB/s | 1.7 TB/s | 80GB | 989 TFLOPS | ||||||
| TSMC-SoW | 3 | 8 | 3.35 TB/s | 1.7 TB/s | 80GB | 989 TFLOPS | ||||||
@@
Metric: We measure pure MoE layer performance during the decode stage as modern LLM serving systems show a trend toward fine-granularity disaggregation. Traditional LLMs benefit from separating prefill and decode stages across different machines, as demonstrated by DistServe [91] and subsequent works [79, 51]. For MoE models, this disaggregation extends further. MegaScale-Infer [92] separates attention and MoE operations onto different machines for optimal batch sizes. Following this trend, we focus on optimizing MoE operations during the decode stage.
Hardware Configuration: We evaluate two multi-chiplet topologies: Tesla Dojo [65, 66] and the TSMC SoW roadmap [70]. As summarized in Table I, Dojo uses a 55 2D mesh, while TSMC SoW adopts an 83 2D mesh. These choices reflect a deployed system (Dojo) and near-future industry support (TSMC SoW).
For both the Dojo and TSMC SoW configurations, each chiplet is H100-like, providing 1,000 TFLOPS FP16 compute, 80GB HBM, 3.35TB/s local HBM bandwidth, and 1.7 TB/s inter-die bandwidth to adjacent chiplets. We also include an extended experiment in Sec. V-F with a Dojo-Enhanced configuration, where each die is B300-like to reflect an anticipated hardware performance trend in the future. We reserve 10% of DRAM for system and hardware management.
Baseline Configurations: We compare our approach against the simple strategy currently used by GPU. The Base configuration adopts an EP-like data placement and assigns an equal number of experts to each die. However, the entire wafer operates as a single large GPU: each die handles the same amount of expert computation without considering expert placement. EP assigns each expert’s computation to the die where it resides, as also adopted by MoEntwine [68]. This eliminates all D2D communication but can cause severe workload imbalance. Note that even under EP, our Global CP and Local CP architecture remains necessary, as expert placement information is still required. We implement three variants: Allo Only uses solely our task allocation strategy; Pred Only includes only the data-driven predictor; and Allo+Pred combines both techniques. These configurations evaluate the individual and combined effects of our proposed methods.
Models and Workloads: We conduct evaluations with real traces collected from Qwen3 and Deepseek V3. The traces are gathered from diverse datasets, including MMLU [25], MMLU Pro [75], ChineseSimpleQA [24], and LiveCodeBench [32], comprising over 24,000 requests per model. Each test batch is filled by sequentially adding requests in the order of MMLU, MMLU-Pro (CH), ChineseSimpleQA, and LiveCodeBench until the target batch size is reached.
@@ V-B Validation of Simulator
We validate our simulator using real measurements from an 8×H100 DGX server. We evaluate both single-GPU execution and two-GPU peer-to-peer (P2P) communication. For single-GPU execution, we benchmark one expert in a MoE layer, which consists of three GEMM operations, across varying batch sizes for both DeepSeek and Qwen. For P2P communication, we measure data migration between two GPUs over payload sizes ranging from 4 KB to 4 GB. To ensure simulation fidelity, we calibrate key parameters to fit the measured data. As shown in Figure 13, the simulator’s error remains within 5% for all test cases.
@@ V-C Throughput
Comparison across models: Our Allo+Pred strategy achieves 7.0, 8.2, 7.3, and 4.1 throughput improvement on Deepseek, Kimi, Llama, and Qwen, respectively. Deepseek and Kimi benefit more due to their larger expert count (256 vs. 128) and more complex selection patterns.
Comparison across chiplet architectures: Our strategy shows 6.0 improvement on Dojo and 7.5 on TSMC, despite similar die counts (25 vs. 24). TSMC’s rectangular layout places dies farther apart, introducing more inter-unit communication without strategic task allocation, hence the larger gain under our strategy.
Comparison with EP: At small batch sizes such as 4096, our strategy and EP perform similarly: few tokens per expert make execution memory-bound, so splitting one expert across multiple dies offers no benefit, and our strategy degenerates to EP. The advantage emerges at larger batches, achieving 1.44 speedup over EP at batch size 16,384.
@@ V-D Hop Reduction
We report hop counts in Figure 12 to show the reduction in inter-unit communication. Hop count is the sum of Manhattan distances for all cross-unit communications. Higher hop counts indicate frequent cross-die data movement. We normalize results to baseline and report hop reduction ratios, where a ratio of 10 means the hop count is reduced to 1/10.
Pred Only reduces hop counts by 4.5, aligning with performance improvement of 3.0. This indicates cross-unit communication is the primary bottleneck in baseline, and reducing hop counts proportionally improves performance.
Allo Only reduces hop counts by 142, exceeding the performance improvement of 6.3. This shows that with our allocation algorithm, inter-unit communication is no longer the sole bottleneck. While reducing hop counts still improves performance, the improvement is not proportional.
Allo+Pred reduces hop counts by over 213 compared to baseline. However, performance improvement is only 6.63 over baseline, with just 1.1 average improvement over Allo Only. This demonstrates that hop count is no longer a performance bottleneck. With the help of our task allocation algorithm, most tasks are distributed to local dies holding related experts, with only extremely popular experts requiring remote allocation. This leads to minimal D2D traffic and shifts the bottleneck to workload distribution.
@@ V-F Comparison with Host CPU-Based Implementation
V-G Area and Power Overhead
We estimate the area and power overhead of all added modules in Table II. Our parameters support up to 100 layers with 512 experts per layer, well beyond the largest existing MoE model (Kimi-K2: 61 layers, 384 experts). The full heatmap for 100 layers totals 50MB and is stored in Global CP DRAM, with an on-chip heatmap cache of 0.5MB (512-bit width) buffering one layer at a time.
The Prediction Table is implemented in registers due to its small size, while all other components use SRAM. Register synthesis is performed with Yosys [77], and SRAM is modeled with CACTI [3], both scaled to 5nm to match the H100 process node. Area estimates for the Global CP and Local CPs are derived from ARM core data. As shown in Table II, the total area and power overhead is less than 0.04%.
| Capacity |
| Prediction Table | 128 B | 16 bit | 25 | 0.0020 | 55.75 | ||||||||
| Address Translation Unit | 4.25 KB | 68 bit | 25 | 0.0048 | 334.25 | ||||||||
| Local CP (A72) [54] | N/A | N/A | 25 | 7.5000 | 7000 | ||||||||
| Expert Distribution Table | 4.5 KB | 72 bit | 1 | 0.0002 | 13.94 | ||||||||
| Heatmap Cache | 0.5 MB | 512 bit | 1 | 0.0278 | 184.67 | ||||||||
| Global CP (A76) [74] | N/A | N/A | 1 | 1.1000 | 1000 | ||||||||
| Total | 6.13 | 8588.61 | |||||||||||
| Overhead (25-die wafer) | 0.04% | 0.04% |
V-H Discussion
@@
insights, which constitute the paper’s primary contribution. Specifically, we follow Insight 3 for task allocation and leverage temporal relation insights (Insight 1 and part of Insight 2) to build a data-driven predictor. However, the remaining insights offer significant potential to further enhance the performance of wafer-scale GPUs in future works. For instance, Insight 2 could inform LLC prefetch strategies and enhance task allocation with LLC data placement awareness; spatial relation insights (Insight 4 and Insight 5) could optimize initial expert placement, further refined by task information from Insight 6. Importantly, our insights extend far beyond this case study to benefit diverse future systems, including multi-GPU clusters [92, 89], CXL-based or CPU-based memory disaggregation [5, 15], Flash-based multi-hierarchy systems [86, 56], and other emerging architectures.
VI Case Study2: Decode Expert Placement with Prefill Information on Current GPUs
VI-A Introduction
Workload imbalance is one of the biggest challenges in large-scale MoE serving (200+ GPUs). EPLB [12] addresses this by dynamically adjusting expert placement, but it triggers every 3000+ steps and relies on periodically collected profiling data [73]. A natural question then arises: how to set expert placement for the initial 1000 decode tokens when no profiling data are yet available? This is especially pressing for short-output requests, for which EPLB never collects enough data to be effective. Inspired by Insight 1, which reveals temporal correlation between prefill and decode stages, we propose leveraging prefill-stage expert selection information to guide expert placement for initial decode steps.
As shown in Figure 16, we design two placement algorithms (details in algorithm 2). The Remap-based algorithm keeps the number of experts per GPU unchanged and reassigns experts across GPUs for a more balanced workload: it sorts experts by decreasing roofline cost and greedily assigns each to the least-loaded GPU, subject to a uniform capacity of experts per GPU. The Duplication-based algorithm reserves extra expert slots on each GPU and uses prefill traces to duplicate hot experts, thereby avoiding congestion: starting from the default contiguous layout (e.g., experts 0–15 on GPU,0, 16–31 on GPU,1, etc.), it greedily adds up to extra replicas per GPU, selecting at each step the (expert, GPU) pair that maximally reduces the bottleneck load ; tokens of a replicated expert are evenly split among all its copies. Both algorithms use a roofline-based cost model to estimate per-GPU load.
VI-B Methodology
We deploy Qwen3-235B with SGLang on 8H100 GPUs with NVLink. We build a distributed profiler by inserting cuda.Event timers into SGLang to measure individual operations (attention, top-, all-to-all, and MoE) on each GPU independently. We manipulate expert placement through SGLang’s init_expert_location interface and use DeepEP as the MoE backend. The ep_dispatch_algorithm is set to “dynamic” so that tokens are evenly distributed across replicas of a duplicated expert.
Metric. We report MoE computation time, i.e., all three expert linear layers, excluding attention, all-to-all, and top-.
Model and Benchmark. We evaluate on Qwen3-235B (94 MoE layers, 128 experts per layer, 8 selected). We use MMLU and Global-MMLU datasets, following the original ordering. Batch sizes range from 64 to 16,384.
Baselines. Default is the standard contiguous placement used by Qwen and SGLang (experts 0–15 on GPU-0, 16–31 on GPU-1, etc.). Best and Worst are the theoretically optimal and worst placements generated with oracle decode-stage selections (not available in practice). Remap and Dup are our two prefill-guided strategies. For Dup, we use one extra slot per GPU, yielding 128+8=136 experts per layer.
VI-C Results
As shown in Figure 17, Remap and Dup achieve speedups of 15.5% and 12.5% over Default, respectively, and deliver over 2 speedup compared with Worst. Both remain within 10% of Best, which exploits decode-stage information unavailable in practice, demonstrating the effectiveness of our approach. Since the two algorithms perform comparably, one can choose between them to fit different memory and system constraints.
We note that our 8-GPU EP scale inherently limits the achievable improvement: with EP8, each GPU holds 16 experts per layer, so every GPU likely contains a mix of hot and cold experts, naturally yielding a relatively balanced workload even under the Default layout (the max/min execution-time ratio is only about 1.3). We expect greater speedups at larger EP scales where load imbalance is more pronounced.
VII Related Work
@@
Several MoE model tech reports [46, 33, 9, 61] provide the MoE routing patterns as part of their evaluation.
For example, the Mixtral report [33] shows the temporal locality of expert selection by reporting the percentage of repetitive assignment.
The OLMoE report [46] shows the co-activation pattern and domain specialization among the experts.
A blog post from SGLang [18] shows expert distribution statistics for the DeepSeekV3 model and the inherent imbalance in expert selection and similarity between prefill and decode.
@@
Various prior work [31, 38, 21, 35, 14, 13, 37, 59, 5, 87, 89, 30] have focused on improving the efficiency of MoE LLM inference by reducing data movement.
For example, Lina and SmartMoE [38, 88] exploit expert selection skewness to dynamically schedule resources during inference and balance traffic across GPUs.
LYNX [21] dynamically reduces active experts while preserving model accuracy.
Pre-gate MoE [31] uses a pre-gating function to alleviate the dynamic nature of expert selection.
Sida [13] builds an offline hash function to predict expert usage and reduce data movement between CPU and GPU.
MoE-Lightning [5] leverages a CPU-GPU pipeline and paged weights to improve resource utilization.
Eliseev and Mazur [14] exploit expert locality and leverage LRU caching to manage GPU and CPU memory.
This work targets data movement reduction in MoE LLM inference. Our cross-model profiling reveals optimization principles that apply broadly to current and future systems regardless of scale.
@@
As single-chip scaling slows, wafer-scale and chiplet packaging offer promising paths for improved compute efficiency. Prior work targets either interconnect design [52, 83, 85, 39, 6, 17, 16] or data placement for specific algorithms [23, 67] and applications [55, 81, 50]. In contrast, we are the first to study MoE LLM serving on wafer-scale GPUs and propose data-movement-centric HW/SW co-design optimizations.
VIII Conclusion
In this work, unlike prior MoE serving studies that adopt system-centric approaches and develop deployment-specific strategies, we study MoE serving from a model-focused perspective. We conduct comprehensive data-movement-centric profiling of the latest large-scale MoE models (200B–1000B) to extract system-agnostic insights. These insights reveal structured patterns underlying seemingly random data movement, providing actionable guidance for designing diverse future MoE serving systems. We validate our findings on both a future wafer-scale GPU architecture and existing multi-GPU systems, achieving significant performance improvements through minimal architectural modifications and lightweight software design, respectively, demonstrating the effectiveness and broad applicability of our insights.